In a sample of 75 workers from a factory in city A, 8% did not graduate from high school. In a sample of 60 workers from a factory in city B, 10% did not graduate from high school. Can we conclude that there is a difference in the propertions of workers who did not graduate from high school in the two cities? Ise a significance level of 0.05.
Answer #1

p1cap = X1/N1 = 6/75 = 0.08
p1cap = X2/N2 = 6/60 = 0.1
pcap = (X1 + X2)/(N1 + N2) = (6+6)/(75+60) = 0.0889

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: p1 = p2
Alternate Hypothesis, Ha: p1 ≠ p2

P-value Approach
P-value = 0.6818
As P-value >= 0.05, fail to reject null hypothesis.

There is not sufficient evidence to conclude that the two proportions are different